# Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation

### v1

@@ Appendix A Base Query Modification
For the ASQA [^7] Apache 2.0 License (Stelmakh et al., 2022) and ExpertQA [^8] MIT License (Malaviya et al., 2023) dataset, we do not modify the base query. For the Longform [^9] MIT License (Köksal et al., 2024) dataset, as there are additional format-related instructions concatenated with the base query (e.g., Respond in 3 sentences.), we eliminate them by using regular expressions. Moreover, we find that Longform dataset contains noisy queries (e.g., This does not provide enough information for an answer to be provided.), which are unfiltered artifacts generated by large language models. In this case, we manually filter similar expressions.
@@ Appendix D Analysis on Intent Operations
We decompose the result of Table 4 according to intent operations in Table 7, focusing on SFT and DPO-QPlanner scores. We discover that C2 queries based on Exclusion score lower than those on the intent of Inclusion. This result aligns with our assumption in Section 3.3.3, where $q_{cov}$ with Exclusion operation require more complicated selection of desirable outline.
@@ Appendix F Training Details
We utilize publicly available software [^10] https://github.com/huggingface/alignment-handbook in our finetuning and alignment stage. We train each stage one epoch using 16 NVIDIA A100 GPUs (40GB of memory). Table 8 describes detailed hyperparameters for each stage.
@@ Appendix G Additional Information on Human Evaluation
We conduct two human evaluation studies separately. For both studies, crowd workers are recruited from Prolific [^11] https://www.prolific.com/. At the beginning of the evaluation, workers are informed what task they are expected to do, there are no foreseeable benefits and risks, their participation is voluntary, and they can leave if they want (see Figure 5).
@@ G.2 Response Evaluation
QPlanner-DPO QPlanner-DPO
Vanilla RAG Unsatisfactory Satisfactory QPlanner-SFT Unsatisfactory Satisfactory
@@
We offer ten single sessions for evaluation (5 sessions for Vanilla RAG vs. QPlanner-DPO and 5 sessions for QPlanner-SFT vs. QPlanner-DPO). Each session has ten evaluators. If wanted, evaluators can participate in more than one session; thirteen out of 63 evaluated multiple sessions. A session takes 20 to 40 minutes. In a session, evaluators are first provided with a short tutorial with evaluation guidelines and examples. Then, they evaluate a pair of responses that answer the same query for twenty queries. They are told that all of the queries are formatted as [question] ($q_{base}$) + [instruction] ($q_{cov}$) and asked to mark a response “Satisfactory” if the response satisfies any of the two evaluation items: (1) the response indeed answers the question [question] with evidence or partial evidence, which includes “there is no evidence but here is useful information,” and (2) the response follows the instruction, [instruction], and mark it “Unsatisfactory” otherwise. If all of the two responses to the same query are rated “Satisfactory,” they were asked to choose which one was a better answer (155 cases in Search Query Evaluation and 392 cases in Response Outline Evaluation; refer to Table 10). We describe a better answer would have more items described above satisfied, or it would be better at following the instruction.
@@ G.2.1 Document Retrieval
For simulating RAG pipeline, we utilize DuckDuckgo [^12] https://serpapi.com/duckduckgo-search-api to search relevant documents. To balance the number of documents, top-10 documents are retrieved in Vanilla RAG, and top-2 documents are retrieved for queries (i.e., 2 * 5 = 10 documents including C2 query) in QPlanner-SFT and QPlanner-DPO. Additionally, we follow the associative selection process, suggested in Lee et al. (2023), to extract relevant evidence paragraph from retrieved documents. Specifically, we construct FLAN-T5-Large (Chung et al., 2024) trained with Wikipedia-based datasets such as MS-MARCO (Bajaj et al., 2018), ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2022), and Qasper (Dasigi et al., 2021). The trained language model matches passages in each document with given queries and returns answerability score deciding whether the paired query and passage are relevant. We select top-1 passage for each document as an evidence for generating RAG response.

### v2

@@ Appendix A Base Query Modification
For the ASQA [^10] Apache 2.0 License (Stelmakh et al., 2022) and ExpertQA [^11] MIT License (Malaviya et al., 2024) dataset, we do not modify the base query. For the Longform [^12] MIT License (Köksal et al., 2024) dataset, as there are additional format-related instructions concatenated with the base query (e.g., Respond in 3 sentences.), we eliminate them by using regular expressions. Moreover, we find that Longform dataset contains noisy queries (e.g., This does not provide enough information for an answer to be provided.), which are unfiltered artifacts generated by large language models. In this case, we manually filter similar expressions.
@@ Appendix D Analysis on Intent Operations
We decompose the result of Table 4 according to intent operations in Table 7, focusing on SFT-QPlanner and DPO-QPlanner scores. We discover that C2 queries based on Exclusion score lower than those on the intent of Inclusion. This result aligns with our assumption in Section 3.3.3, where $q_{cov}$ with Exclusion operation require more complicated selection of desirable outline.
@@ Appendix F Training Details
We utilize publicly available software [^13] https://github.com/huggingface/alignment-handbook in our finetuning and alignment stage. We train each stage one epoch using 16 NVIDIA A100 GPUs (40GB of memory). Table 8 indicates detailed hyperparameters for each stage.
@@ Appendix G Additional Information on Human Evaluation
We conduct two human evaluation studies separately (i.e., outline and RAG evaluation). For both studies, crowdworkers are recruited from Prolific [^14] https://www.prolific.com/. At the beginning of the evaluation, workers are informed what task they are expected to do, there are no foreseeable benefits and risks, their participation is voluntary, and they can leave if they want (see Figure 5).
@@ G.2 RAG Response Evaluation
DPO-QPlanner DPO-QPlanner
Vanilla RAG Unsatisfactory Satisfactory SFT-QPlanner Unsatisfactory Satisfactory
@@
We offer ten single sessions for evaluation (5 sessions for Vanilla RAG vs. DPO-QPlanner and 5 sessions for SFT-QPlanner vs. DPO-QPlanner). Each session has ten evaluators. If wanted, evaluators can participate in more than one session; thirteen out of 63 evaluated multiple sessions. A session takes 20 to 40 minutes. In a session, evaluators are first provided with a short tutorial with evaluation guidelines and examples. Then, they evaluate a pair of responses that answer the same query for twenty queries. They are told that all of the queries are formatted as [question] ($q_{base}$) + [instruction] ($q_{cov}$) and asked to mark a response “Satisfactory” if the response satisfies any of the two evaluation items: (1) the response indeed answers the question [question] with evidence or partial evidence, which includes “there is no evidence but here is useful information,” and (2) the response follows the instruction, [instruction], and mark it “Unsatisfactory” otherwise. If all of the two responses to the same query are rated “Satisfactory,” they were asked to choose which one was a better answer (155 cases in Search Query Evaluation and 392 cases in Response Outline Evaluation; refer to Table 10). We describe a better answer would have more items described above satisfied, or it would be better at following the instruction.
@@ G.2.1 Document Retrieval
For simulating RAG pipeline, we utilize DuckDuckgo [^15] https://serpapi.com/duckduckgo-search-api to search relevant documents. To balance the number of documents, top-10 documents are retrieved in Vanilla RAG, and top-2 documents are retrieved for each subquery (i.e., 2 * 5 = 10 documents including C2 query) in SFT-QPlanner and DPO-QPlanner. Additionally, we follow the associative selection process, suggested in Lee et al. (2023), to extract relevant evidence paragraphs from retrieved documents. Specifically, we construct FLAN-T5-Large (Chung et al., 2024) trained with Wikipedia-based datasets such as MS-MARCO (Bajaj et al., 2018), ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2022), and Qasper (Dasigi et al., 2021). The trained language model matches passages in each document with given subqueries and returns an answerability score deciding whether the paired subquery and passage are relevant. We select the top-1 passage for each document as evidence for generating RAG response.